OpenStudy (anonymous):

y=4-3x finding the inverse function. PLEASE show steps

OpenStudy (anonymous):

Is it multiplying by the x or dividing?

OpenStudy (amriju):

y=(4-x)/3...see basically express the function in the form x=f(y)...now put x in place of y and y in place of x

OpenStudy (anonymous):

Either way I don't think it matters much, the first you would do is switch the position of x and y giving you x=4-3y, you want to isolate the y so you have to subtract 4 from sides giving x-4=-3y then you just divide by -3 giving you y=-1/3x+4/3

OpenStudy (anonymous):

seems like multiplying @JoshR1213

OpenStudy (anonymous):

Thank you so much @JoshR1213 @amriju

OpenStudy (amriju):

welcome :)